OpenStudy (anonymous):

May someone please help me understand how to work this problem out? x - 3 1/4 = 5 6/8

OpenStudy (texaschic101):

do you know how to turn mixed numbers into improper fractions ?

OpenStudy (texaschic101):

3 1/4 = (3 x 4 + 1)/4 = 13/4 5 6/8 = (5 x 8 + 6)/8 = 46/8 x - 3 1/4 = 5 6/8 -- turn to improper fractions x - 13/4 = 46/8 -- add 13/4 to both sides to get x by itself x - 13/4 + 13/4 = 46/8 + 13/4 -- combine like terms x = 46/8 + 13/4 -- common denominator is 8 x = 46/8 + 26/8 x = 72/8 which reduces to 9